OpenStudy (anonymous):

C(52, 5) = a. 5,804,600 b. 2,598,960 c. 3,385,900 d. 10,922,600

OpenStudy (anonymous):

\[ \frac{52!}{5! (52-5)!}=2598960 \] b

OpenStudy (anonymous):

\[ \frac{52!}{5! (52-5)!}=\frac{52!}{5! 47!}=\frac{48\times49\times 50\times\times 51 \times 52}{120}=2598960 \]